Integrate Sved Vädx with n=4 using (1) Trapezoidal rule and (2) Simpsons rule, and compare with the true value.
Answer #1

Solution 8 Trade Te ftphone= (1.4 3 [235]* = { [ 8%] = { 1.22.627 = 15-0849 3/2 3/2 Ia Io 15.0849 True value / Exact value Iy = f(n)=√x Yo = f(0) 2 Ýi = f/2) = 12 1.414 Y2 = f (4) = 54 Y3-316) = 6 = 2.449 44 = 818) - 2.828 8 11: [19o+yn)+2(91+42 + 9(11) [2= Suda uring simpsons sule Here, we will consider simpsons Ird scele [ha 2 2020 2122 22-4 x 22=6 19-8 Yooo 411.414 Y